OpenStudy (anonymous):

Last year, Kiri's age was a multiple of 7. This year, Kiri's age is a multiple of 5. In how many years will Kiri be 26 years old? (STEP BY STEP ANSWER PLEASE!)

OpenStudy (dumbcow):

11 years 7k = Age -1 5m = Age 7k = 5m -1 try m=1,2,3... m=3 7k =15-1 = 14 5*3=15 26-15 = 11

OpenStudy (anonymous):

Oh I get it now! Thanks :D